OpenStudy (anonymous):

what is n if √n=√(2n-6)?

OpenStudy (ash2326):

proceed as earlier square both sides take 2n to left side and then find n

OpenStudy (anonymous):

I got n=2n-6 n+2n=6 n+n=3 What did I do wrong?

OpenStudy (unklerhaukus):

n=2n-6 (minus two n) -n=-6 (divide by negative 1) n=6

OpenStudy (anonymous):

Thank you :)